A woman has $1.70 in dimes and nickels. She has 10 more nickels than dimes. How many dimes does she have?
Answer by mananth(16949)

let dimes be x
nickels will be x+10
..
value 170 cents
...
10x+5(x+10)=170
10x+5x+50=170
15x=120
/15
x= 8 dimes
nickels = 18
...
8*10+18*5=170
...
m.ananth@hotmail.ca